# EchoMimicV3: 1.3B Parameters are All You Need for Unified Multi-Modal and Multi-Task Human Animation

### v3

@@ Abstract
Figure 1: EchoMimicV3, an efficient framework with just 1.3B parameters for unified Multi-Modal and Multi-Task human video generation. Based on the audio, prompt and reference image over various scenes, EchoMimicV3 generate a video where lip movements and facial expression align with the audio, and the gestures and scene follow the prompt.
@@
Codehttps://github.com/antgroup/echomimic˙v3
Introduction
@@
Recent advancements in human animation have been boosted by large-scale video diffusion models (LVDM) (Wang et al. 2025; Chen et al. 2025a; Yang et al. 2024; Wan et al. 2025; Kong et al. 2024; Wei et al. 2025; Cui et al. 2024; Lin et al. 2025). While the adoption of LVDMs has led to higher quality and stronger generalization in human animation, it also introduces notable challenges: 1) prohibitive training costs and slow inference speeds due to LVDM’s massive parameter scale; 2) complex model routing caused by the need for separate expert LVDMs for different animation tasks (e.g., text-to-video, image-to-video, lip-sync). Unfortunately, the latter challenge further aggravates the inefficiencies of the former.
@@
Soup-of-Tasks for Multi-Task Unification. We reformulate diverse animation tasks from a novel spatiotemporal reconstruction perspective akin to Masked Autoencoders (MAE) (Kong et al. 2025). Specifically, lip-syncing is recast as mouth spatial region reconstruction, while Text-to-Video (T2V), Image-to-Video (I2V), First-Last-Frame-to-Video (FLF2V) can be viewed as intermediate temporal frames reconstruction. EchoMimicV3 unifies these tasks by fully exploiting the common masked sequence input inherent in mainstream video diffusion models. By doing so, step-wise diffusion model and patch-wise reconstruction elegantly converge without painful trade-offs.
@@
Inspired by the PhD Loss in EchoMimicV2 (Meng et al. 2025), our motivation of the Multi-Modal PhDA stems from the observation that different modals exhibit varying levels of importance across different timestep phases. Specifically, text conditions maintain consistent importance throughout phases, image conditions are most influential during the early and middle timestep phases, and audio conditions are particularly relevant in the initial phase. The Multi-Modal PhD allocates weights to each modal expert branch based on this Phase-specific Modal Importance Law, fusing them through linear combination.
Novel Training and Inference Strategy. We propose a novel training strategy to effectively integrate the aforementioned Soup-of-Tasks and Soup-of-Modals paradigms. Traditional post-training methods like Direct Preference Optimization (DPO) (Wallace et al. 2024) reject undesired distributions with preference data but suffer from high computational costs, limited generalization, and sensitivity to preference data quality. We propose Negative DPO, which uses pairing-free negative samples to reject the distribution of preference negative data. By interleaving Negative DPO with Supervised Fine-Tuning (SFT), we dynamically mitigate spatial inconsistencies (e.g., identity preservation issues) and temporal artifacts (e.g., color shifts) in Negative DPO-SFT cycle training. Notably, our Negative DPO simplifies the conventional DPO pipeline while achieving effective performance.
@@
EchoMimicV3
Overview
Pipeline. In this section, we present EchoMimicV3, a unified framework for multi-task and multi-modal human animation, as shown in Fig. 2. EchoMimicV3 generates talking human videos with conditions of the reference image, audio, and text prompt, without cumbersome conditions such as predefined 2D or 3D poses. To unify diverse tasks, we utilize a new Soup-of-Task paradigm (Sec 3.2). Additionally, to enhance the multi-modal capabilities of lightweight models with 1.3B parameters, we propose a novel Soup-of-Modal paradigm (Sec 3.3). Furthermore, to ensure the stability of our framework containing various tasks and multiple modals, we propose a training strategy, in which a new Negative DPO is injected into SFT process for dynamically rejecting undesirable results. Correspondingly, we introduce a Phase-aware Negative CFG (PNG) to enhance the rejection of negative results for inference (Sec 3.4).
Soup-of-Tasks
Spatial-temporal masked reconstruction. In the transformer of Wan-Fun series video model (Wallace et al. 2024), a corresponding 0-1 masked sequence is concatenated with the video latent as the input. Inspired by MAE (He et al. 2022), we view diverse tasks from the spatial-temporal reconstruction perspective: diverse tasks can be viewed as employing distinct masking strategies, with variations exit solely in the mask inputs. Specifically, the mask input for the audio-driven Text-to-video (T2V), Image-to-Video (I2V), First-Last-Frame-to-Video (FLF2V), and lip syncing are designed as $\mathsf{M}_{T2V}$, $\mathsf{M}_{I2V}$, $\mathsf{M}_{FLF2V}$ and $\mathsf{M}_{lip}$, respectively, as illustrated in Fig. 2, where each task is encoded with a unique 0-1 sequence pattern. This design facilitates the integration of multiple tasks into a unified model without necessitating architectural modification.
Soup-of-Tasks training strategy includes a counter-intuitive task schedule and implicit task mixture approach. Intuitively, curriculum learning involves progressing from easier (task with highest mask ratio) to more challenging tasks (task with lower mask ratio). However, in contrast, we adopt a counter-intuitive task schedule: we begin with the most difficult tasks with the highest masking ratio and gradually incorporate other simpler tasks. The reason for doing so is mainly that the challenging tasks align more closely with the task paradigm of the pretrained model.
@@
Multi-Modal PhDA is designed for allocating the $\mathbb{W}(c,\tau)$ for different modals, according to the timestep $\tau$ in each iteration. The weights are sampled from a meta distribution, based on the Phase-specific Modal Importance inspired by by the PhD Loss (Meng et al. 2025). The weights for different modals in different timestep phases can be calculated as follows:
@@
Training Strategy
Negative Direct Preference Optimization (Negative DPO) leverages suboptimal negative samples generated from intermediate checkpoints of the Supervised Fine-Tuning (SFT) to iteratively refine the model. Specifically, during SFT, denote intermediate checkpoints as $\{\mathcal{M}_{\tau}\mid\tau\in I\}$, where $\tau$ is the number of iterations. We sample checkpoints $\{\mathcal{M}_{s_{i}}\mid{s_{i}}\in I\}$, where $s_{i}$ denotes the number of iterations for the $i$ training stage. Note that each training stage has the same issue. For the $i$ training stage, we employ $\mathcal{M}_{s_{i}}$ as our reference model to generate videos $\mathcal{D}_{s_{i}}$, then we identity negative samples from $\mathcal{D}_{s_{i}}$, and then annotate the issue based on user feedback, omitting the pairing of $\langle p,y^{-},y^{+}\rangle$. The negative preference data for stage $i$ can be denoted as $\langle p^{-},y^{-}\rangle\in\mathcal{D}_{s_{i}}$.
@@
Dataset. We train our model on a combination of the EchoMimicV2 dataset, the HDTF dataset (Zhang et al. 2021), and additional self-collected data. To ensure data quality, we apply preprocessing steps, including audio synchronization and subtitle removal. The total training data comprises approximately 1,500 hours of video content.
Metrics. To comprehensively evaluate our model’s performance, we employ the following metrics: 1) Image Quality: Fréchet Inception Distance (FID) (Deng et al. 2019) is used to assess the fidelity and visual quality of generated images. 2) Video Quality: Fréchet Video Distance (FVD) (Unterthiner et al. 2018) measures temporal coherence and overall video quality. 3) Perceptual and Aesthetic Analysis: We analyze perceptual quality (IQA) and aesthetic appeal (ASE) of the generated content. 4) Audio-Visual Alignment: For lip-sync tasks, we use Sync-C and Sync-D metrics (Prajwal et al. 2020) to evaluate synchronization accuracy. 5) Consistency Metrics: We adopt Vbench2.0 (Zheng et al. 2025) metrics, including Identity Consistency (ID), Human Anatomy (HA), Clothing Consistency (HC), and Background Consistency (BC). For quantitative evaluation, we randomly select 300 videos generated by EchoMimicV3.
Results
Comparison with SOTA Methods. We conducted both qualitative and quantitative comparisons with existing methods for half-body digital human generation. These include EchoMimicV2 (Meng et al. 2025), HunyuanAvatar (Chen et al. 2025b), OmniAvatar (Gan et al. 2025), Hallo3 (Cui et al. 2024), and MultiTalk (Sung-Bin et al. 2024). As shown in Table 1, our method is competitive in various evaluation metrics, such as audio-lip synchronization, human motion accuracy, identity preservation, video aesthetics, and overall video quality, even compared to methods with $10\times$ parameters (e.g., FantasyTalk). Notably, we achieve superior performance in identity preservation, video aesthetics, background consistency, clothing fidelity, and body motion precision over competing methods.
The normalized results per dimension for clearer comparison are illustrated in Fig. 3. Additionally, we conduct a user study to validate our method, as shown in Fig. 3. Notably, our approach achieves superior Chinese lip-sync accuracy and human motion fidelity compared to state-of-the-art methods. Furthermore, a qualitative comparison with SOTA methods is provided in Fig. 4 to demonstrate the effectiveness of our method.
@@ Results
Ablation on Efficiency. Our method achieves talking head generation within 5 inference steps. It can be seen in Table 1, compared to 14B-parameter models such as FantasyTalk and HunyuanAvatar, EchoMimicV3 provides a $18\times$ speedup. For talking human synthesis, leveraging TeaCache (Liu et al. 2025), our approach generates a 5-second video in approximately 4 minutes with 25 inference steps.
Related Work
Video Generation
Early research in video generation focused on 2D U-Net architectures and motion modules, later extended to 3D frameworks. The incorporation of spatiotemporal convolutions enabled models like SVD to produce short clips with high spatial fidelity. Diffusion-based models, such as DiT, further advanced the field by leveraging self-attention mechanisms, reducing inductive bias, and improving temporal coherence and motion fluidity. Recent progress in text-to-video synthesis has been propelled by frameworks like CogVideo (Yang et al. 2024), which combine large-scale vision-language pretraining with hierarchical token fusion for precise semantic alignment. HunyuanVideo (Kong et al. 2024) introduced a dual-stream DiT architecture, enhancing conditional control in complex scenarios. Wan (Wan et al. 2025) demonstrated scalable models with state-of-the-art performance, setting new benchmarks and highlighting the potential for broader video synthesis applications.
Human Animation
The objective of audio-driven human animation is to synthesize natural, expressive gestures from speech, ensuring semantic, emotional, and temporal alignment. While prior work has focused on animating talking heads (Zhu et al. 2024; Xu et al. 2024), recent methods have advanced quality and consistency. For instance, EMO (Tian et al. 2024) uses a Frame Encoding module for temporal coherence, and AniPortrait (Wei, Yang, and Wang 2024) maps 3D facial structures to 2D poses for coherent sequences. Vlogger (Zhuang et al. 2024) generates high-fidelity videos using a diffusion-based framework. EchoMimic (Chen et al. 2024) supports flexible generation modes, driven by audio, facial poses, or both. CyberHost (Lin et al. 2024) enables versatile animations with multimodal controls, including hand poses and body trajectories. OmniHuman (Lin et al. 2025) achieves state-of-the-art performance using progressive dropout strategies. Hunyuanvideo-Avatar (Chen et al. 2025b) and MultiTalk (Sung-Bin et al. 2024) support multi-audio-driven multi-character animations, while Omniavatar (Gan et al. 2025) injects audio efficiently into transformers. However, DiT-based methods still struggle with lip sync and long-duration video generation, highlighting the need for further advancements.
Conclusion

### v4

@@ Abstract
1 Introduction
@@
Recent advancements in human animation have been boosted by large-scale video diffusion models (LVDM) [20, 2, 24, 19, 8, 21, 4, 11]. While the adoption of LVDMs has led to higher quality and stronger generalization in human animation, it also introduces notable challenges: 1) prohibitive training costs and slow inference speeds due to LVDM’s massive parameter scale; 2) complex model routing caused by the need for separate expert LVDMs for different animation tasks (e.g., text-to-video, image-to-video, lip-sync). Unfortunately, the latter challenge further aggravates the inefficiencies of the former.
@@
Soup-of-Tasks for Multi-Task Unification. We reformulate diverse animation tasks from a novel spatiotemporal reconstruction perspective akin to Masked Autoencoders (MAE) [9]. Specifically, lip-syncing is recast as mouth spatial region reconstruction, while Text-to-Video (T2V), Image-to-Video (I2V), and First-Last-Frame-to-Video (FLF2V) can be viewed as intermediate temporal frames reconstruction. EchoMimicV3 unifies these tasks by fully exploiting the common masked sequence input inherent in mainstream video diffusion models. By doing so, step-wise diffusion model and patch-wise reconstruction elegantly converge without painful trade-offs.
@@
Inspired by the PhD Loss in EchoMimicV2 [13], our motivation of the Multi-Modal PhDA stems from the observation that different modals exhibit varying levels of importance across different timestep phases. Specifically, text conditions maintain consistent importance throughout phases, image conditions are most influential during the early and middle timestep phases, and audio conditions are particularly relevant in the initial phase. The Multi-Modal PhD allocates weights to each modal expert branch based on this Phase-specific Modal Importance Law, fusing them through linear combination.
Novel Training and Inference Strategy. We propose a novel training strategy to effectively integrate the aforementioned Soup-of-Tasks and Soup-of-Modals paradigms. Traditional post-training methods like Direct Preference Optimization (DPO) [18] reject undesired distributions with preference data but suffer from high computational costs, limited generalization, and sensitivity to preference data quality. We propose Negative DPO, which uses pairing-free negative samples to reject the distribution of preference negative data. By interleaving Negative DPO with Supervised Fine-Tuning (SFT), we dynamically mitigate spatial inconsistencies (e.g., identity preservation issues) and temporal artifacts (e.g., color shifts) in Negative DPO-SFT cycle training. Notably, our Negative DPO simplifies the conventional DPO pipeline while achieving effective performance.
@@
2 EchoMimicV3
2.1 Overview
Pipeline. In this section, we present EchoMimicV3, a unified framework for multi-task and multi-modal human animation, as shown in Fig. 1. EchoMimicV3 generates talking human videos with conditions of the reference image, audio, and text prompt, without cumbersome conditions such as predefined 2D or 3D poses. To unify diverse tasks, we utilize a new Soup-of-Task paradigm (Sec 3.2). Additionally, to enhance the multi-modal capabilities of lightweight models with 1.3B parameters, we propose a novel Soup-of-Modal paradigm (Sec 3.3). Furthermore, to ensure the stability of our framework containing various tasks and multiple modals, we propose a training strategy, in which a new Negative DPO is injected into the SFT process for dynamically rejecting undesirable results. Correspondingly, we introduce a Phase-aware Negative CFG (PNG) to enhance the rejection of negative results for inference (Sec 3.4).
2.2 Soup-of-Tasks
Spatial-temporal masked reconstruction. In the transformer of Wan-Fun series video model [18], a corresponding 0-1 masked sequence is concatenated with the video latent as the input. Inspired by MAE [7], we view diverse tasks from the spatial-temporal reconstruction perspective: diverse tasks can be viewed as employing distinct masking strategies, with variations existing solely in the mask inputs. Specifically, the mask input for the audio-driven Text-to-video (T2V), Image-to-Video (I2V), First-Last-Frame-to-Video (FLF2V), and lip syncing are designed as $\mathsf{M}_{T2V}$, $\mathsf{M}_{I2V}$, $\mathsf{M}_{FLF2V}$ and $\mathsf{M}_{lip}$, respectively, as illustrated in Fig. 1, where each task is encoded with a unique 0-1 sequence pattern. This design facilitates the integration of multiple tasks into a unified model without necessitating architectural modification.
Soup-of-Tasks training strategy includes a counter-intuitive task schedule and implicit task mixture approach. Intuitively, curriculum learning involves progressing from easier (task with the highest mask ratio) to more challenging tasks (task with the lower mask ratio). However, in contrast, we adopt a counter-intuitive task schedule: we begin with the most difficult tasks with the highest masking ratio and gradually incorporate other simpler tasks. The reason for doing so is mainly that the challenging tasks align more closely with the task paradigm of the pretrained model.
@@ 2.3 Soup-of-Modals
Multi-Modal PhDA is designed for allocating the $\mathbb{W}(c,\tau)$ for different modals, according to the timestep $\tau$ in each iteration. The weights are sampled from a meta distribution, based on the Phase-specific Modal Importance inspired by the PhD Loss [13]. The weights for different modals in different timestep phases can be calculated as follows:
@@
2.4 Training Strategy
Negative Direct Preference Optimization (Negative DPO) leverages suboptimal negative samples generated from intermediate checkpoints of the Supervised Fine-Tuning (SFT) to iteratively refine the model. Specifically, during SFT, denote intermediate checkpoints as $\{\mathcal{M}_{\tau}\mid\tau\in I\}$, where $\tau$ is the number of iterations. We sample checkpoints $\{\mathcal{M}_{s_{i}}\mid{s_{i}}\in I\}$, where $s_{i}$ denotes the number of iterations for the $i$ training stage. Note that each training stage has the same issue. For the $i$ training stage, we employ $\mathcal{M}_{s_{i}}$ as our reference model to generate videos $\mathcal{D}_{s_{i}}$, then we identify negative samples from $\mathcal{D}_{s_{i}}$, and then annotate the issue based on user feedback, omitting the pairing of $\langle p,y^{-},y^{+}\rangle$. The negative preference data for stage $i$ can be denoted as $\langle p^{-},y^{-}\rangle\in\mathcal{D}_{s_{i}}$.
@@ 3.1 Experimental Setups
Dataset. We train our model on a combination of the EchoMimicV2 dataset, the HDTF dataset [25], and additional self-collected data. To ensure data quality, we apply preprocessing steps, including audio synchronization and subtitle removal. The total training data comprises approximately 1,500 hours of video content.
Metrics. To comprehensively evaluate our model’s performance, we employ the following metrics: 1) Image Quality: Fréchet Inception Distance (FID) [5] is used to assess the fidelity and visual quality of generated images. 2) Video Quality: Fréchet Video Distance (FVD) [17] measures temporal coherence and overall video quality. 3) Perceptual and Aesthetic Analysis: We analyze perceptual quality (IQA) and aesthetic appeal (ASE) of the generated content. 4) Audio-Visual Alignment: For lip-sync tasks, we use Sync-C and Sync-D metrics [14] to evaluate synchronization accuracy. 5) Consistency Metrics: We adopt Vbench2.0 [26] metrics, including Identity Consistency (ID), Human Anatomy (HA), Clothing Consistency (HC), and Background Consistency (BC). For quantitative evaluation, we randomly select 300 videos generated by EchoMimicV3.
3.2 Results
Comparison with SOTA Methods. We conducted both qualitative and quantitative comparisons with existing methods for half-body digital human generation. These include EchoMimicV2 [13], HunyuanAvatar [1], OmniAvatar [6], Hallo3 [4], and MultiTalk [15]. As shown in Table 1, our method is competitive in various evaluation metrics, such as audio-lip synchronization, human motion accuracy, identity preservation, video aesthetics, and overall video quality, even compared to methods with $10\times$ parameters (e.g., FantasyTalk). Notably, we achieve superior performance in identity preservation, video aesthetics, background consistency, clothing fidelity, and body motion precision over competing methods.
The normalized results for each dimension are illustrated in Fig. 2, and the corresponding user study validation is also shown within the same figure. Notably, our approach achieves superior Chinese lip-sync accuracy and human motion fidelity compared to state-of-the-art methods. Furthermore, a qualitative comparison with SOTA methods is provided in Fig. 3 to demonstrate the effectiveness of our method.
@@ 3.3 Ablation Studies
Ablation on Efficiency. Our method achieves talking head generation within 5 inference steps. It can be seen in Table 1, compared to 14B-parameter models such as FantasyTalk and HunyuanAvatar, EchoMimicV3 provides a $18\times$ speedup. For talking human synthesis, leveraging TeaCache [12], our approach generates a 5-second video in approximately 4 minutes with 25 inference steps.
4 Related Work
4.1 Video Generation
Early research in video generation focused on 2D U-Net architectures and motion modules, later extended to 3D frameworks. The incorporation of spatiotemporal convolutions enabled models like SVD to produce short clips with high spatial fidelity. Diffusion-based models, such as DiT, further advanced the field by leveraging self-attention mechanisms, reducing inductive bias, and improving temporal coherence and motion fluidity. Recent progress in text-to-video synthesis has been propelled by frameworks like CogVideo [24], which combine large-scale vision-language pretraining with hierarchical token fusion for precise semantic alignment. HunyuanVideo [8] introduced a dual-stream DiT architecture, enhancing conditional control in complex scenarios. Wan [19] demonstrated scalable models with state-of-the-art performance, setting new benchmarks and highlighting the potential for broader video synthesis applications.
4.2 Human Animation
The objective of audio-driven human animation is to synthesize natural, expressive gestures from speech, ensuring semantic, emotional, and temporal alignment. While prior work has focused on animating talking heads [27, 23], recent methods have advanced quality and consistency. For instance, EMO [16] uses a Frame Encoding module for temporal coherence, and AniPortrait [22] maps 3D facial structures to 2D poses for coherent sequences. Vlogger [28] generates high-fidelity videos using a diffusion-based framework. EchoMimic [3] supports flexible generation modes, driven by audio, facial poses, or both. CyberHost [10] enables versatile animations with multimodal controls, including hand poses and body trajectories. OmniHuman [11] achieves state-of-the-art performance using progressive dropout strategies. Hunyuanvideo-Avatar [1] and MultiTalk [15] support multi-audio-driven multi-character animations, while Omniavatar [6] injects audio efficiently into transformers. However, DiT-based methods still struggle with lip sync and long-duration video generation, highlighting the need for further advancements.
5 Conclusion